OpenStudy (anonymous):

In the system shown below, what are the coordinates of the solution that lies in quadrant IV? x^2+y^2=25 x-y^2=-5

OpenStudy (anonymous):

x2−y2=25 (x+y)(x−y)=25 x+y=25 x−y=1 add to get 2x=26 x=13 then solve for y

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

answer is (13,12) (4,3) does not work because 4 + 3 = 7, not 25

OpenStudy (anonymous):

If you solve the system by elimination, we note that adding both equations together gives you \(\large x^2+x = 20\implies x^2+x-20=0\). This equation can be factored, so we see that \(\large x^2+x-20=-\implies (x+5)(x-4)=0\). Therefore, \(\large x=-5\) or \(\large x=4\). In order for the point of intersection to be in the fourth quadrant, the x value needs to be positive, but the y value needs to be negative. Therefore, if \(\large x=4\), plugging this into the first equation gives us \(\large 4^2+y^2 = 25\implies y^2 = 9 \implies y=\pm 3\). But as I mentioned before, we need y to be negative in order to get the point in the fourth quadrant. Hence, the solution that we want is \(\large (4,-3)\). I've attached a graph of the two equations, so you can clearly see that this point is in fact the solution we're after. I hope this makes sense! :-)